On the problem Log (base 1/2)|6x-9|= Log (base 1/2)|-6| , what Logarithm rules do I use? And how do I got about implementing them into this problem to get the answer? I don't even know where to start lol, Thanks in advance.
Answer by lwsshak3(6517)

Log (base 1/2)|6x-9|= Log (base 1/2)|-6|
**
Since the bases are the same the numbers must also equal to each other, that is:
|6x-9|=|-6|
|6x-9|=6
solving for two conditions
6x-9=6
6x=15
x=15/6=5/2
..
-(6x-9=6
-6x+9=6
6x=9-6=3
x=3/6=1/2
ans:
x=5/2
or
x=1/2